OpenStudy (anonymous):

can please help me with math question if f(x)=x^2+1 and g(x)=2x, find [fog](x)

OpenStudy (anonymous):

f of g of x means you take x, plug it into g(x)... then whatever that gives you, plug that into f(x) So in this case you take x, plug it into g(x) and get 2x Now take 2x and plug that into f(x) ... so you get (2x)^2 + 1

OpenStudy (anonymous):

okay my teacher told me the answer was 4x^2+1 i need to know how so i can keep up in class

OpenStudy (anonymous):

(2x)^2 + 1 = 2^2 * x^2 + 1 = 4x^2 + 1

OpenStudy (anonymous):